### v2

# Instruct-MusicGen: Unlocking Text-to-Music Editing for Music Language Models via Instruction Tuning

Abstract
Recent advances in text-to-music editing, which employ text queries to modify music (e.g. by changing its style or adjusting instrumental components), present unique challenges and opportunities for AI-assisted music creation. Previous approaches in this domain have been constrained by the necessity to train specific editing models from scratch, which is both resource-intensive and inefficient; other research uses large language models to predict edited music, resulting in imprecise audio reconstruction. To Combine the strengths and address these limitations, we introduce Instruct-MusicGen, a novel approach that finetunes a pretrained MusicGen model to efficiently follow editing instructions such as adding, removing, or separating stems. Our approach involves a modification of the original MusicGen architecture by incorporating a text fusion module and an audio fusion module, which allow the model to process instruction texts and audio inputs concurrently and yield the desired edited music. Remarkably, Instruct-MusicGen only introduces $\sim$8new parameters to the original MusicGen model and only trains for 5K steps, yet it achieves superior performance across all tasks compared to existing baselines, and demonstrates performance comparable to the models trained for specific tasks. This advancement not only enhances the efficiency of text-to-music editing but also broadens the applicability of music language models in dynamic music production environments.  [^1] Work done during Yixiao’s internship at Sony AI. [^2] Code, model weights and demo are available at: https://bit.ly/instruct-musicgen.
1 Introduction
The rapid advances in text-to-music generation have opened up new possibilities for AI-assisted music creation (Agostinelli et al., 2023; Copet et al., 2023; Li et al., 2023; Liu et al., 2023a; Chen et al., 2023). This paradigm shift has also sparked a growing interest in developing models that offer greater controllability (Lin et al., 2023; Wu et al., 2023a; Melechovsky et al., 2023; Lin et al., 2024) and editability (Han et al., 2023; Zhang et al., 2024; Hussain et al., 2023) over the music generation process. In music production, a stem—a mixed group of tracks often related by instrument type (like drums or lead vocals)—is essential for mixing and mastering because it allows producers to isolate, adjust, and manipulate individual elements of a song. Following the definition in  Zhang et al. (2024), "text-to-music editing" involves using textual queries to modify various aspects of a music recording, which can be categorised into two main types: intra-stem editing, which focuses on modifying a single stem (e.g., changing the instrument, timbre, or performance style), and inter-stem editing, which involves altering the relationships among stems (e.g., adding, removing, or separating stems). Our work mainly focuses on the problem of inter-stem editing.
Previous attempts to develop text-based music editing models have encountered several challenges. Some approaches (Wang et al., 2023; Han et al., 2023) have focused on training specialized editing models from scratch, which is resource-intensive and may not yield results comparable to state-of-the-art music generation models. Other work (Zhang et al., 2023b; Hussain et al., 2023; Yang et al., 2023) has sought to leverage existing language models (LLMs) and MusicGen (Copet et al., 2023), allowing the LLM to interpret editing instructions without further training the music model. Although this approach offers flexibility, it often lacks the ability to precisely reconstruct the conditional audio, leading to unreliable results. To address these limitations, an ideal solution should harness the knowledge embedded in pretrained models to ensure high-quality audio output while adapting the architecture to accommodate the specific requirements of music editing tasks.
Figure 1: Comparison between MusicGen and instruct-MusicGen. Instruct-MusicGen accepts both audio input and editing instruction text as conditions.
In this paper, we introduce Instruct-MusicGen, a novel approach that applies an instruction-following tuning strategy to the pretrained MusicGen model, enhancing its ability to follow editing instructions effectively without finetuning all its parameters. As shown in Figure 1, by incorporating an audio fusion module based on LLaMA-adapter (Zhang et al., 2023a; Lin et al., 2023) and a text fusion module based on LoRA (Hu et al., 2021) into the original MusicGen architecture, we allow the model to process both precise audio conditions and text-based instructions simultaneously, while the original MusicGen does not. This enables Instruct-MusicGen to perform a wide range of editing tasks, such as adding, separating, and removing stems. To train Instruct-MusicGen, we synthesize an instructional dataset using the Slakh2100 dataset (Manilow et al., 2019), introducing only  8% additional parameters compared to the original model, and finetune the model for only 5K steps, which is less than 1% of training a music editing model from scratch.
We evaluate Instruct-MusicGen on two datasets: the Slakh test set and the out-of-domain MoisesDB dataset (Pereira et al., 2023). Our model outperforms existing baselines and achieves performance comparable to models specifically trained for individual tasks. This demonstrates the effectiveness of our approach in leveraging pretrained models for text-to-music editing while maintaining high-quality results. The main contributions of this work are threefold:
  1. 1.
    We propose Instruct-MusicGen, a novel approach that finetunes pretrained large music language models to follow editing instructions effectively. This method addresses the limitations of previous approaches by leveraging pretrained models to ensure high-quality audio output and precise reconstruction, combining both music and text multi-modal control.
  2. 2.
    We enhance the capabilities of a pretrained text-to-music model to perform multiple tasks, including adding, separating, and extracting stems from music audio, all within a single training process. This approach significantly reduces the computing resources cost compared to training specialised editing models from scratch.
  3. 3.
    We extend the benchmarking for text-to-music editing tasks, providing a comprehensive comparison across various models. Our model outperforms existing baselines and achieves performance comparable to models specifically trained for individual tasks, demonstrating the effectiveness and versatility of our approach.
2 Related work
Despite the growing body of research on text-to-music generation (Agostinelli et al., 2023; Copet et al., 2023; Li et al., 2023; Yao et al., 2023; Melechovsky et al., 2023; Evans et al., 2024a, b; Chen et al., 2023), the creation of music remains an inherently iterative process, involving multiple rounds of editing and revision. This ongoing need for refinement and adjustment underscores the necessity for specialized models that facilitate and enhance music editing.
Text-based music editing provides a flexible approach for editing music using textual queries. This method is similar to other modalities that require editing, such as image (Brooks et al., 2023; Liu et al., 2023b) and video (Chai et al., 2023; Ceylan et al., 2023) editing. In text-to-music editing, tasks leverage text to make precise alterations to existing music compositions. To address these needs in music editing, AUDIT (Wang et al., 2023) and InstructME (Han et al., 2023) developed a diffusion model trained with paired music editing data. Additionally, models like M2UGen Hussain et al. (2023), Loop Copilot (Zhang et al., 2023b), MusicAgent (Yu et al., 2023), ComposerX (Deng et al., 2024) and WavCraft (Liang et al., 2024) use large language models (LLMs) for reasoning and regenerate music with external music generation models. Furthermore, Postolache et al. (2024) attempt to jointly model multi-stem distribution of music for text-based generation and separation. Certain models focus exclusively on specific tasks within music editing, such as conditioned generation (Lin et al., 2023, 2024) and separation (Liu et al., 2023c), along with intra-stem editing tasks like text-based timbre transfer and style transfer (Zhang et al., 2024; Manor and Michaeli, 2024; Li et al., 2024).
The task of inter-stem music editing is closely related to stem-wise music generation, which is also crucial in the music production process. While these efforts are not directly tied to text-based controls, some research focuses on modeling stem-wise representations to enable simultaneous stem generation and separation. For instance, Jen-1 Composer (Yao et al., 2023) and MSDM (Mariani et al., 2023) jointly model the distribution of music with four stems using a diffusion model. The abilities of most existing stem-wise music models are restricted to a fixed set of 4 stems, which limits flexibility but enhances controllability; In contrast, StemGen (Parker et al., 2023) introduces a novel approach by employing a non-autoregressive Transformer architecture for more flexible stem-wise generation, moving beyond the constraints of a fixed stem set.
Our work distinguishes itself from these existing efforts in several key ways. First, rather than developing a new model from scratch or strictly adhering to a fixed set of stems, we leverage the power of a pretrained music language model, MusicGen, and enhance it with instruction tuning. This approach not only reduces the computational cost but also retains the high audio quality of the original MusicGen model. By integrating both audio and text fusion modules, our model, Instruct-MusicGen, can precisely interpret and execute complex editing instructions involving multiple stems. Furthermore, our method introduces minimal additional parameters and requires significantly less training, demonstrating a more efficient and scalable solution for text-based music editing. This dual approach of using pretrained models with specialized adapters sets our work apart, providing a flexible and powerful tool for the evolving needs of AI-assisted music creation.
3 Method
3.1 MusicGen
The original MusicGen consists of three components: (1) the EnCodec (Défossez et al., 2022) audio encoder and decoder, which compress music audio waveforms into latent codes and reconstruct them back into waveforms; (2) a multi-layer transformer architecture that models sequences of latent codes, capturing higher-level music representations and efficiently modeling internal relationships within music audio; and (3) the T5 (Raffel et al., 2020) text encoder, which converts text descriptions into embeddings for text-conditioned generation.
MusicGen uses EnCodec (Défossez et al., 2022), a convolutional auto-encoder, to map real audio waveforms to latent codes. EnCodec employs Residual Vector Quantization (RVQ) (Zeghidour et al., 2022) to compress audio into tokens using multiple codebooks, where each quantizer encodes the quantization error from the previous one. For a reference audio $X\in\mathbb{R}^{d\cdot f_{n}}$, where $d$ is the duration and $f_{n}$ is the sample rate, EnCodec compresses $X$ into $Q\in\{1,...,L\}^{N\times d\cdot f_{s}}$, where $L$ is the RVQ codebook size, $N$ is the number of codebooks, and $f_{s}$ is the latent code sample rate ($f_{s}<<f_{n}$). In MusicGen, $N=4$, $f_{n}=50$, $f_{s}=32000$, and $L=2048$. Finally, The transformer models the sequence relationships over latent codes.
3.2 Instruct-MusicGen
MusicGen is a text-to-music generation model, capable of generating music audio from a given text prompt. However, MusicGen cannot edit existing music audio. To address this limitation, we introduce Instruct-MusicGen, which transforms MusicGen into a model that can follow editing instructions to modify existing music audio.
Instruct-MusicGen takes a music audio input $X_{\text{cond}}$ and a text instruction $X_{\text{instruct}}$ (e.g., "Add guitar") as inputs. The model then edits the music audio $X_{\text{cond}}$ according to the instruction $X_{\text{instruct}}$ and generates the desired edited music $X_{\text{music}}$. As illustrated in Figure 2, Instruct-MusicGen incorporates two additional modules into the vanilla MusicGen: an audio fusion and a text fusion module.
Figure 2: The architecture of instruct-MusicGen. The audio fusion module transforms the conditional music audio into embeddings using a duplicated encoder and integrates these embeddings into the MusicGen decoder. The text fusion module modifies the cross-attention mechanism to handle text instructions by finetuning specific layers while keeping the text encoder parameters frozen.
3.2.1 Audio Fusion Module
The audio fusion module enables Instruct-MusicGen to accept external audio inputs, which is inspired by Zhang et al. (2023a) and Lin et al. (2023). Figure 2 presents the diagram of the audio fusion module. Initially, we convert $X_{\text{cond}}$ into EnCodec tokens, followed by re-encoding these tokens into the embedding $z_{\text{cond}},z_{\text{cond}}$ through the pre-trained embedding layers of MusicGen. Similarly, we transform $X_{\text{music}}$ into the pretrained embedding $z_{\text{music}}$.
The module begins with duplicating self-attention modules of the pretrained MusicGen model to extract latent representations of $z_{\text{cond}}$. Given that MusicGen consists of $M$ layers, we denote $Z_{\text{cond}}=\{z^{0}_{\text{cond}},z_{\text{cond}}^{1},\ldots,z_{\text{cond}}^{M}\}$ and $Z_{\text{music}}=\{z^{0}_{\text{music}},z_{\text{music}}^{1},\ldots,z_{\text{music}}^{M}\}$ to represent the hidden states of $z_{\text{cond}}$ and $z_{\text{music}}$ respectively. Note that we use a learnable input embedding as $z^{0}_{\text{cond}}$ and initialize $z^{0}_{\text{music}}$ with $z_{\text{music}}$.
Let $W_{\rm Q}$, $W_{\rm K}$, $W_{\rm V}$ and $W_{\rm O}$ denote the query, key, value, and output projection matrices of the pretrained self-attention transformer layer respectively, then, $z_{\text{music}}$ is projected into the latent space,
$K=Z_{\text{music}}W_{\text{K}},\hskip 14.22636ptQ=Z_{\text{music}}W_{\text{Q}},\hskip 14.22636ptV=Z_{\text{music}}W_{\text{V}}.$ (1)
The attention output $O$ of $Q$, $K$, and $V$ is computed as follows,
$O={\rm Attention}(Q,K,V)W_{\rm O},$ (2)
${\rm Attention}(Q,K,V)=\text{Softmax}\left(\frac{QK^{\mathsf{T}}}{\sqrt{d}}\right)V,$ (3)
where $d$ is the dimension of $K$.
For $z_{\text{cond}}$, we design a learnable linear layer $\text{Linear}_{\text{cond}}$ to incorporate information from $X_{\text{cond}}$ into MusicGen, and $M$ layer-wise learnable position embeddings $\{e_{1},e_{2},....e_{M}\}$ to enhance the sequital modeling. Let $Z^{\prime}_{\text{cond}}=\{z_{\text{cond}}^{1}+\text{Linear}_{\text{cond}}(z_{\text{cond}})+e_{1},z_{\text{cond}}^{2}+\text{Linear}_{\text{cond}}(z_{\text{cond}})+e_{2},\ldots,z_{\text{cond}}^{M}+\text{Linear}_{\text{cond}}(z_{\text{cond}})+e_{M}\}$, then, $Z^{\prime}_{\text{cond}}$ is projected into the latent space,
$K^{\prime}=Z^{\prime}_{\text{cond}}W_{\text{K}},\hskip 14.22636ptQ^{\prime}=Z^{\prime}_{\text{cond}}W_{\text{Q}},\hskip 14.22636ptV^{\prime}=Z^{\prime}_{\text{cond}}W_{\text{V}}.$ (4)
We compute the attention output of $Q^{\prime}$, $K^{\prime}$, and $V^{\prime}$ and pass it through each layer sequentially. Let ${Q^{\prime}}_{m}$, ${K^{\prime}}_{m}$, and ${V^{\prime}}_{m}$ denote $m$-th layer value in $Q^{\prime}$, $K^{\prime}$, and $V^{\prime}$, which are computed from $z_{\text{cond}}^{m-1}$, then,
$z_{\text{cond}}^{m}=\text{Attention}({Q^{\prime}}_{m},{K^{\prime}}_{m},{V^{\prime}}_{m})W^{m}_{\text{O}}.$ (5)
To fuse information of $Z_{\text{cond}}$ into $Z_{\text{music}}$, we compute the cross attention between $Z_{\text{cond}}$ and $Z_{\text{music}}$,
${\rm CrossAttention}(Q,K^{\prime},V^{\prime})=\text{Softmax}\left(\frac{(Q+Q^{\prime})(K^{\prime})^{\mathsf{T}}}{\sqrt{d}}\right)V^{\prime}.$ (6)
Finally, the attention output of $Z_{\text{music}}$ is updated as follows,
$O^{\prime\prime}={\rm CrossAttention}(Q,K^{\prime},V^{\prime})W_{\rm O},$ (7)
$O_{\text{fuse}}=O+g\cdot O^{\prime\prime},$ (8)
where $g\in\mathbb{R}$ is a zero-initialized learnable gating factor.
Thus, the total trainable parameters in Instruct-MusicGen include the input embedding $z^{0}_{\rm cond}$, linear layers ${\rm Linear}_{\rm cond}$, learnable position embeddings $\{e_{1},e_{2},....e_{M}\}$, and learnable gating factors $g$.
3.2.2 Text Fusion Module
To replace the text description input with instruction input, we modify the behavior of the current text encoder. We achieve this by finetuning only the cross-attention module between the text embedding and the music representations while keeping the text encoder’s parameters frozen.
First, the text is embedded and encoded by the T5 text encoder as $z_{\text{instruct}}=\text{T5}(X_{\text{instruct}})$. Let $W^{\prime}_{\rm Q}$, $W^{\prime}_{\rm K}$, and $W^{\prime}_{\rm V}$ denote the query, key, and value projection matrices of the pretrained cross-attention transformer layers respectively. Then, the query, key, and value of $z_{\text{instruct}}$ and $Z_{\text{music}}$ for the multi-modal cross-attention are calculated as follows:
$K_{\rm instruct}=z_{\text{instruct}}W^{\prime}_{\rm K},\hskip 14.22636ptQ_{\rm instruct}=Z_{\text{music}}W^{\prime}_{\rm Q},\hskip 14.22636ptV_{\rm instruct}=z_{\text{instruct}}W^{\prime}_{\rm V}.$ (9)
For efficient finetuning of the cross-attention module, we apply LoRA to the projection layers $W^{\prime}_{Q}$ and $W^{\prime}_{V}$. LoRA decomposes each of $W^{\prime}_{Q}$ and $W^{\prime}_{V}$ into two smaller low-rank matrices, significantly reducing the number of parameters requiring fine-tuning. The $Q_{\rm instruct}$ and $V_{\rm instruct}$ of cross-attention are then revised as follows:
$Q_{\rm instruct}=z_{\text{instruct}}(W^{\prime}_{\rm Q}+(W_{\rm Q}^{\text{A}})^{\mathsf{T}}W_{\rm Q}^{\text{B}}),$ (10)
$V_{\rm instruct}=z_{\text{instruct}}(W^{\prime}_{\rm V}+(W_{\rm V}^{\text{A}})^{\mathsf{T}}W_{\rm V}^{\text{B}}).$ (11)
During fine-tuning, $W^{\prime}_{Q}$, $W^{\prime}_{K}$, and $W^{\prime}_{V}$ are frozen, while $W^{\text{A}}_{Q}$, $W^{\text{B}}_{Q}$, $W^{\text{A}}_{V}$, and $W^{\text{B}}_{V}$ are trainable.
4 Experiments
4.1 Dataset
For our objective evaluations, we utilise two distinct datasets, each serving a specific purpose in assessing both in-domain and out-of-domain performance capabilities of various models.
  1. 1.
    Slakh2100 dataset (Manilow et al., 2019). The Synthesized Lakh (Slakh) Dataset, originally derived from the Lakh MIDI Dataset v0.1, comprises high-quality audio tracks synthesised using sample-based virtual instruments. Known as Slakh2100, this dataset features 2100 tracks complete with corresponding MIDI files. These tracks are methodically divided into subsets designated for training (1500 tracks), validation (375 tracks), and testing (225 tracks), collectively amounting to approximately 145 hours of audio mixtures. This dataset primarily supports audio source separation tasks within a controlled, synthesised environment.
  2. 2.
    MoisesDB dataset (Pereira et al., 2023). In contrast to the Slakh2100, the MoisesDB dataset caters to source separation with a broader scope. It includes 240 real audio tracks sourced from 45 diverse artists spanning twelve musical genres. Uniquely, MoisesDB organises its tracks into a detailed two-level hierarchical taxonomy of stems, offering a varied number of stems per track, each annotated with textual descriptions. This setup facilitates a more comprehensive evaluation across different genres and stem configurations.
The rationale for selecting two datasets lies in their diverse configurations and common applications. While the Slakh dataset is traditionally utilized for training models tailored to a four-stem arrangement (e.g., Postolache et al., 2024), our model, Instruct-MusicGen, although initially trained on this dataset, is designed to generalise to various stem configurations. Conversely, models such as InstructME and AUDIT are trained on private or larger, more diverse datasets. By employing both Slakh2100 and MoisesDB, we ensure a comprehensive evaluation, allowing us to fairly compare the adaptability and performance of different models under varying conditions of data familiarity and complexity.
4.2 Data Preprocessing
We utilized the Slakh2100 dataset to construct an instruction-based dataset for our experiments, employing the following pipeline:
  1. 1.
    A data point was randomly selected from the Slakh training dataset.
  2. 2.
    An instruction was chosen from a predefined set {add, remove, extract} along with a target stem. Subsequently, $n$ other stems were selected from the remaining stems.
  3. 3.
    An offset was randomly determined to cut a 5-second audio clip. If the target stem contained more than 50% silence, a different offset was selected.
  4. 4.
    The stems were mixed according to the specified instructions to create a triplet consisting of {instruction text, condition audio input, audio ground truth}.
4.3 Experimental setup
For the finetuning of MusicGen, we jointly trained the audio fusion module and the text fusion module. The optimisation process utilised the AdamW optimiser, with a learning rate set at $5e^{-3}$. We use L2 loss over latent tokens as the training objective. Training incorporated a Cosine Annealing scheduler with an initial warmup of 100 steps. The training regimen extended over 5,000 steps with an accumulated batch size of 32, achieved through setting the batch size to 8 and using gradient accumulation over 4 iterations. The finetuning process was executed on a single NVIDIA A100 GPU and was completed within a span of two days.
4.4 Baselines
In this section, we explore various baseline models, each distinguished by their unique methodologies for handling audio data. We categorise these into three types: text-based music editing models, stem-wise modelling, and specialised models for single text-based tasks (see Table 1 for a comparison). Note we use "ME" to distinguish whether a model is a text-to-music editing model or not.
4.4.1 Text-based Music Editing Models
  1. 1.
    AUDIT (Wang et al., 2023): AUDIT is an instruction-guided audio editing model, consisting of a variational autoencoder (VAE) for converting input audio into a latent space representation, a T5 text encoder for processing edit instructions, and a diffusion network that performs the actual audio editing in the latent space. The system accepts mel-spectrograms of input audio and edit instructions, and generates the edited audio as output.
  2. 2.
    InstructME (Han et al., 2023): InstructME is a novel framework for instruction-guided music editing and remixing that leverages latent diffusion models. It enhances a U-Net with multi-scale aggregation to ensure consistency in the music before and after editing, and incorporates a chord progression matrix into the semantic space to improve melodic harmony. To handle extended musical pieces, InstructME utilizes a chunk transformer to recognize long-term temporal dependencies within music sequences.
  3. 3.
    M2UGen (Hussain et al., 2023): The M2UGen framework leverages large language models to comprehend and generate music across various modalities, integrating abilities from external models such as MusicGen (Copet et al., 2023) and AudioLDM 2 (Liu et al., 2023a). It is designed to stimulate creative outputs from diverse sources, showcasing robust performance in multi-modal music generation.
4.4.2 Specialised models
  1. 1.
    GMSDI (Postolache et al., 2024): GMSDI generalizes multi-source diffusion models to handle compositional music generation tasks without requiring pre-separated data. It allows for the flexible parameterization of an arbitrary number of sources and employs text embeddings to enhance semantic control over the music generation process. GMSDI uses the Slakh and Jamendo datasets for model training.
  2. 2.
    AudioSep (Liu et al., 2023c): AudioSep introduces the LASS-Net, which performs audio source separation based on natural language queries. This model effectively combines acoustic and linguistic data to isolate audio sources that match the language descriptions, significantly improving over conventional methods. AudioSep uses multiple datasets including AudioSet, VGGSound, AudioCaps, Clotho v2 and WavCaps dataset for model training.
Model ME Param size Dataset Hours (h) Training steps
AUDIT 942M (1.5B) Multiple $\sim$6500 0.5M
InstructME 967M (1.7B) Multiple 417 2M
M2UGen 637M ($\sim$9B) MUEdit 60.22 -
GMSDI $\sim$1B Slakh 145 $\sim$1M
AudioSep 238M Multiple $\sim$14000 1M
Ours 264M (3.5B) Slakh 145 5K
Table 1: Comparison of different models.
4.5 Metrics
To comprehensively evaluate the quality and alignment of predicted music with ground truth and textual descriptions, we employ several metrics. The Fr’echet Audio Distance (FAD)(Kilgour et al., 2019) [^3] https://github.com/gudgud96/frechet-audio-distance. measures overall audio quality by comparing the distributions of feature embeddings between two sets of audio files. The CLAP score (CLAP)(Wu et al., 2023b) [^4] https://github.com/LAION-AI/CLAP. assesses how well the audio content aligns with corresponding textual descriptions, indicating semantic coherence. Kullback-Leibler Divergence (KL) [^5] https://github.com/haoheliu/audioldm_eval. evaluates the information loss by comparing the probability distributions of audio features from different sources, ensuring feature similarity between predicted and ground truth music. The Structural Similarity (SSIM)(Wang et al., 2004) metric, adapted from image quality assessment, evaluates the structural similarity between predicted music and the ground truth. Lastly, the Scale-Invariant Signal-to-Distortion Ratio (SI-SDR)(Roux et al., 2019) and Scale-Invariant Signal-to-Distortion Ratio improvement (SI-SDRi) (Isik et al., 2016) measure audio quality, particularly in source separation tasks, by quantifying the distortion relative to a reference signal and the improvement after processing, respectively. A detailed explaination can be found in Appendix A.
4.6 Results
Our evaluation of Instruct-MusicGen demonstrates its superior performance across various tasks compared to existing text-to-music editing baselines (AUDIT, InstructME, M2UGen) and specialized trained models (GMSDI, AudioSep). On the Slakh dataset (Table 2), Instruct-MusicGen excelled in adding, removing, and extracting stems, achieving the lowest Fréchet Audio Distance (FAD) and highest CLAP and SSIM scores in the addition task. It also significantly improved the signal-to-noise ratio (SI-SDR) in the removal task, showing balanced performance across all metrics and proving its robustness in various editing scenarios.  [^6] Since InstructME’s model weights and evaluation protocol are not public released, we only record the evaluation results in Table 5, which were reported in the original paper.
[…]
Add AUDIT 6.88 0.12 1.02 0.21 - -
InstructME* - - - - - -
M2UGen 7.24 0.22 0.99 0.20 - -
GMSDI 3.29 0.17 0.46 0.31 - -
Ours 3.75 0.23 0.67 0.26 - -
Remove AUDIT 15.48 0.07 2.75 0.35 -45.60 -47.28
InstructME* - - - - - -
M2UGen 8.26 0.09 1.59 0.23 -44.20 -46.13
GMSDI 1.30 0.12 0.23 0.78 5.67 3.74
AudioSep 0.90 0.18 0.63 0.65 1.38 -0.31
Ours 3.35 0.12 0.66 0.45 -2.09 -3.77
Extract AUDIT 15.08 0.06 2.38 0.42 -52.90 -50.16
InstructME* - - - - - -
M2UGen 8.14 0.11 2.15 0.31 -46.38 -43.53
GMSDI 2.91 0.12 0.28 0.74 2.73 5.58
AudioSep 2.49 0.15 1.00 0.62 -0.96 1.78
Ours 3.24 0.12 0.54 0.52 -9.00 -6.15
Table 2: Comparison of text-based music editing models on the Slakh dataset (4 stems).
Similarly, in the MoisesDB dataset evaluations (Table 3), Instruct-MusicGen continued to demonstrate strong performance. It achieved competitive FAD scores and showed improvements in CLAP and SSIM metrics for both addition and removal tasks. While it did not always lead in SI-SDR, it consistently outperformed baseline models and was comparable to specialized models, highlighting its efficiency and effectiveness in text-to-music editing applications.
[…]
Add AUDIT 4.06 0.12 0.84 0.21 - -
InstructME* - - - - - -
M2UGen 5.00 0.18 0.83 0.20 - -
GMSDI 5.40 0.12 0.60 0.30 - -
Ours 3.79 0.18 0.35 0.35 - -
Remove AUDIT 10.72 0.10 2.46 0.34 -44.32 -57.10
InstructME* - - - - - -
M2UGen 3.75 0.13 1.27 0.19 -43.94 -56.73
GMSDI 1.42 0.09 0.41 0.65 -2.35 -15.13
AudioSep 1.25 0.13 0.49 0.68 4.97 -7.81
Ours 5.05 0.10 0.84 0.34 -13.70 -26.48
Extract AUDIT 6.67 0.07 1.97 0.45 -54.53 -56.17
InstructME* - - - - - -
M2UGen 5.74 0.08 1.91 0.25 -42.84 -44.49
GMSDI 4.95 0.08 1.32 0.54 -8.19 -9.83
AudioSep 1.81 0.14 1.01 0.65 1.14 -0.50
Ours 4.96 0.11 1.36 0.40 -21.39 -23.03
Table 3: Comparison of text-based music editing models on the MoisesDB dataset.
Despite these promising results, there are limitations to our approach. One significant limitation is the reliance on synthetic data from the Slakh2100 dataset for training, which may not fully capture the diversity and complexity of real-world music recordings. For example, while Instruct-MusicGen achieved a competitive FAD score of 3.35 and an SSIM of 0.45 in the removal task on the Slakh dataset, its SI-SDR improvement (SI-SDRi) was -3.77, indicating that while the audio quality is high, the separation quality could be enhanced. Similarly, in the extraction task, our model achieved a respectable FAD of 3.24 and SSIM of 0.52, but the SI-SDR and SI-SDRi metrics were -9.00 and -6.15, respectively, highlighting challenges in accurately isolating stems without introducing noise. These metrics suggest that while Instruct-MusicGen performs well in preserving audio fidelity and similarity, it struggles with signal clarity in more complex editing scenarios.
We provide further ablation studies in the Appendix A.  [^7] Demos are available at: https://bit.ly/instruct-musicgen.
5 Conclusion
In this paper, we introduced Instruct-MusicGen, a novel approach to text-to-music editing that fosters joint musical and textual controls. By finetuning the existing MusicGen model with instruction tuning, Instruct-MusicGen demonstrated its capability of editing music in various ways, including adding, separating and extracting a stem from music audio using textual queries, without the need for training specialized models from scratch. Also, it outperforms various baseline models that are dedicated to specific music editing tasks. Furthermore, our method uses significantly fewer resources than previous models, with a requirement of tuning only 8% of the parameters of the original MusicGen.
Instruct-MusicGen still has several limitations. Firstly, the generative editing tasks do not guarantee signal-level precision, as the process of predicting EnCodec tokens inherently involves some inaccuracy. Second, the finetuning process still relies on paired data, limiting the model’s applicability to more complex editing tasks where such data is not readily available. These constraints highlight the need for further research to improve signal accuracy and expand the model’s capabilities to handle a broader range of editing scenarios without heavily relying on paired datasets. Nonetheless, the ability to edit music with fewer computational resources opens up new avenues for creative expression and automation in music composition. Instruct-MusicGen represents a significant contribution to the field of text-to-music research, offering a robust and efficient tool for both musicians and researchers.
Appendix A Appendix / supplemental material
A.1 Metrics
The metrics to evaluate model performance are listed below.
  1. 1.
    Fréchet Audio Distance (FAD) [Kilgour et al., 2019] [^8] https://github.com/gudgud96/frechet-audio-distance. measures the similarity between two sets of audio files by comparing multivariate Gaussian distributions fitted to feature embeddings from the audio data. We use the FAD score to evaluate the overall audio quality of the predicted music.
  2. 2.
    CLAP score (CLAP) [Wu et al., 2023b] [^9] https://github.com/LAION-AI/CLAP. evaluates the alignment and relevance of audio content with corresponding textual descriptions, using models pretrained on both audio and language tasks to assess semantic coherence between audio and text. A high CLAP score indicates the predicted music aligns with the text description well.
  3. 3.
    Kullback-Leibler Divergence (KL) [^10] https://github.com/haoheliu/audioldm_eval. assesses the difference between the probability distributions of audio features from two sources, indicating information loss when approximating one distribution with another. A low KL score indicates the predicted music shares similar features with the ground truth.
  4. 4.
    Structural Similarity (SSIM) [Wang et al., 2004] is an image quality metric that we adapt to evaluate structural similarity between predicted music and ground truth.
  5. 5.
    Scale-Invariant Signal-to-Distortion Ratio (SI-SDR) [Roux et al., 2019] quantifies audio quality, especially in source separation tasks. It is scale-invariant, useful for varying audio volumes, and measures distortion relative to a reference signal. We use SI-SDR score to evaluate the signal loss of the predicted audio.
  6. 6.
    Scale-Invariant Signal-to-Distortion Ratio improvement (SI-SDRi) [Isik et al., 2016] extends SI-SDR, measuring the improvement in signal-to-distortion ratio before and after processing. It is commonly used in audio enhancement and separation contexts.
A.2 Ablations
A.2.1 Hyperparameter selection
The total trainable parameters include the input embedding layer, linear layers for joint embeddings learning, and gate scalars, and the linear layers has over 80% parameters in the total parameters. In this ablation we want to investigate whether the parameters can be less by revising the architecture of the linear layers.
The original linear layers has the shape of $(48,2048,2048)$, where 48 is the layer number of MusicGen model, and the $(2048,2048)$ is a linear transformation from embedding to layer-wise representation. We replace a single layer with an auto-encoder architecture, consisting of two linear layers: $(2048,M)$ and $(M,2048)$, where $M\ll 2048$.
We record and compare the loss on the validation set between the variant and the original instruct-MusicGen model, shown in Figure 3. The variation has $M=12$, which reduces the parameter size from 201M to 23M. From the experimental results, we find that even the paramater size becomes less, the instruction tuning still works for MusicGen, but the small parameter size limits the model’s representation ability, making the model performs worse than the original version.
Figure 3: The comparison between the original instruct-MusicGen and its variant with less parameters.
A.2.2 Comparison with original text encoder
We use an audio fusion module to enable MusicGen to accept additional audio input, allowing the model to process instruction text input without adding a new adapter. By finetuning the cross-attention module between the text encoder and the MusicGen transformer, we can fuse instructional knowledge into text information. In this ablation study, we examine the necessity of finetuning the text encoder. Since the text encoder is already capable of accepting text input, can it directly understand the semantic meaning of editing instructions?
To investigate this, we revised the architecture of Instruct-MusicGen’s text fusion module, configuring the music decoder to only accept cross-attention information from the original text encoder.
From Table 4, we observe that removing the text fusion module and relying on the original text encoder to interpret the instructions significantly decreases the model’s performance. Qualitative results from the predicted samples indicate that the original text encoder, which is designed for text description, struggles with the format of instructional commands. It fails to fully understand the semantics of keywords such as "extract," "remove," "no," and "only," which are not typically used to describe music but rather to indicate editing operations. Therefore, the presence of the text fusion module is crucial for MusicGen to correctly interpret the semantic meaning of instructions.
[…]
Add w/o. text fusion 5.26 0.15 0.70 0.25 - -
w. text fusion 3.75 0.23 0.67 0.26 - -
Remove w/o. text fusion 5.58 0.09 1.23 0.31 -10.32 -12.25
w. text fusion 3.35 0.12 0.66 0.45 -2.09 -3.77
Extract w/o. text fusion 7.58 0.04 1.87 0.42 -23.06 -20.21
w. text fusion 3.24 0.12 0.54 0.52 -9.00 -6.15
Table 4: Performance comparison between models with and without text fusion module across different tasks.
A.3 Experiments on other datasets
Since GMSDI only handle 4-stem music modelling, we construct a variant evaluation on Slakh dataset, where all stem can be considered, and the results are shown in Table 5.
[…]
Add AUDIT 6.30 0.08 1.27 0.14 - -
InstructME* 3.15 - - - - -
M2UGen 5.80 0.21 0.88 0.17 - -
Ours 3.07 0.20 0.53 0.21 - -
Remove AUDIT 12.19 0.06 2.58 0.27 -45.40 -49.32
InstructME* 1.87 - - - - -
M2UGen 6.42 0.10 1.42 0.22 -42.72 -46.02
AudioSep 0.96 0.16 0.57 0.64 2.51 -0.78
Ours 2.95 0.12 0.64 0.39 -3.01 -6.31
Extract AUDIT 13.78 0.04 2.86 0.43 -54.95 -49.10
InstructME* 5.04 - - - - -
M2UGen 6.29 0.12 1.64 0.30 -45.65 -42.53
AudioSep 2.81 0.16 1.30 0.59 -2.22 0.91
Ours 3.25 0.08 0.85 0.45 -9.89 -6.77
Table 5: Comparison of Text-Based Music Editing Models on the Slakh dataset (full stems).
A.4 Social Impact
The introduction of text-based music editing models like Instruct-MusicGen has the potential to significantly impact the music industry by democratizing music creation and editing. By lowering the technical barriers to music editing, these models enable a broader range of individuals, including those without formal music training, to engage in creative processes. This can foster greater diversity in music production, as more people from various backgrounds can contribute their unique perspectives and ideas. Additionally, the efficiency and accessibility of text-based music editing can streamline workflows for professional musicians and producers, allowing them to experiment with different musical elements quickly and with less effort.
However, it is important to recognise the limitations and challenges associated with these advancements. The reliance on paired data for finetuning can restrict the model’s versatility in handling more complex editing tasks. Moreover, the inherent loss in generative editing might affect the precision of the output, potentially requiring additional finetuning by human experts. While these models offer new tools for creative expression, they are not a replacement for the nuanced understanding and expertise of skilled musicians. The social impact of text-based music editing models will largely depend on how they are integrated into existing workflows and their acceptance by the music community.
A.5 Audio samples
We provide audio samples in 4. From the results we can find the instruct-MusicGen performs the adding , removing and extracting tasks.
(a) Input music.
(b) Edited music output.
(c) Ground truth.
(d) Input music.
(e) Edited music output.
(f) Ground truth.
(g) Input music.
(h) Edited music output.
(i) Ground truth.
Figure 4: Audio samples. Fig (a-c): Extracting a stem. Instruction: "Extract drum."; Fig (d-f): Adding a stem. Instruction: "Add piano."; Fig (g-i): Removing a stem. Instruction: "Remove bass";